# Null projections and noncommutative function theory in operator algebras

### v1

@@ 5.2. Hilbert function spaces and their multipliers
we infer that ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$ is equivalent to $\Theta({\mathcal{A}}({\mathcal{H}})q)=\Theta({\mathcal{T}}({\mathcal{H}})q)$, or
$\{\varphi:M_{\varphi}\in{\mathcal{A}}({\mathcal{H}})\}\chi_{E}=\mathrm{C}({\mathbb{S}}_{d})\chi_{E}.\qed$
A bounded linear functional $\varphi:{\mathcal{A}}({\mathcal{H}})\to{\mathbb{C}}$ will be said to be ${\mathcal{M}}({\mathcal{H}})$-Henkin if $\varphi$ extends weak-*** continuously to ${\mathcal{M}}({\mathcal{H}})$. We note here that this coincides with the notion used in [CT2023Henkin] by Equation (6) above Lemma 5.6 there (or [DH2023, Lemma 2.1]). This will be used implicitly below, whenever results from [CT2023Henkin] regarding Henkin functionals are invoked.
@@
[…]
A closed projection $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ is ${\mathcal{M}}({\mathcal{H}})$-totally null if $|\varphi|(q)=0$ for every functional $\varphi\in{\mathcal{T}}({\mathcal{H}})^{*}$ such that $\varphi|_{{\mathcal{A}}({\mathcal{H}})}$ is ${\mathcal{M}}({\mathcal{H}})$-Henkin. A Borel subset $E\subset{\mathbb{S}}_{d}$ is said to be ${\mathcal{M}}({\mathcal{H}})$-totally null if $|\mu|(E)=0$ for every ${\mathcal{M}}(H)$-Henkin measure. We clarify the relation between these notions in the next result. In what follows, we will repeatedly use the fact that $\varphi(\cdot(1-z_{K}))|_{{\mathcal{A}}({\mathcal{H}})}$ is an ${\mathcal{M}}({\mathcal{H}})$-Henkin functional for every $\varphi\in{\mathcal{T}}({\mathcal{H}})^{*}$ (see [CT2023Henkin, Lemma 5.5]).
@@
Since $q$ is closed and ${\mathcal{T}}({\mathcal{H}})$ is finitely generated by the monomials, hence is separable, we may choose a sequence of positive contractions $(t_{n})$ in ${\mathcal{T}}({\mathcal{H}})$ decreasing to $q$ in the weak-*** topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. In particular, we see that the bounded sequence of continuous functions $((\tau\circ\gamma)(t_{n}))$ converges pointwise on ${\mathbb{S}}_{d}$ to $\chi_{E}$. By dominated convergence and using that $(\tau\circ\nu)^{**}(z_{K})=1$, we then find
@@ 5.3. The Davidson–Hartz theorems
The previous result has bearing on the topic of peak-interpolation. Indeed, it allows us to provide alternative proofs (and even some refinements) of some of the results from [DH2023] via the noncommutative peak-interpolation theory surveyed in [blecher2013]. For the most part, the statement below is a commutative one. That our approach takes us through the abstract noncommutative theory is a novel feature that we feel has potential for further future applications. We thank Ken Davidson and Michael Hartz for several enjoyable and helpful discussions on this topic.
@@
Let $g\in\mathrm{C}(E)$ with $\|g\|=1$. Let $k_{0}\in\mathrm{C}({\mathbb{S}}_{d})$ be a strictly positive function such that $k_{0}|_{E}=1$ and $10<k<$ 1 on ${\mathbb{S}}_{d}\setminus E$. Let $k\in\mathrm{C}({\mathbb{S}}_{d})$ be strictly positive such that $|g|\leq k$ on $E$ and put $k_{1}=\min\{k,k_{0}\}$. By Tietze’s extension theorem, there is $g_{0}\in\mathrm{C}({\mathbb{S}}_{d})$ such that $g_{0}|_{E}=g$ and $\|g_{0}\|=1$. Recall that $\tau\circ\gamma:{\mathcal{T}}({\mathcal{H}})\to\mathrm{C}({\mathbb{S}}_{d})$ is a surjective ***-homomorphism. Thus, there is $b\in{\mathcal{T}}({\mathcal{H}})$ with $\|b\|=1$ and a positive element $d\in{\mathcal{T}}({\mathcal{H}})$ with $\|d\|=\|k_{1}\|$ such that $(\tau\circ\gamma)(b)=g_{0}$ and and $(\tau\circ\gamma)(d)=k_{1}$.
@@
Using that $E$ is an ${\mathcal{A}}({\mathcal{H}})$-interpolation set, we find ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$ by Lemma 5.4. Since ${\mathcal{T}}({\mathcal{H}})$ is nuclear, by Proposition 4.8 we infer that there is $\kappa>0$ such that $q$ is a complete $({\mathcal{A}}({\mathcal{H}}),\kappa)$-interpolation projection.
@@
The following result closes the circle of results about peak and null sets and projections for ${\mathcal{A}}({\mathcal{H}})$. Recall that by [CT2023Henkin, Lemma 5.5], there is a weak-*** homeomorphic ***-isomorphism $\Omega:{\mathcal{T}}({\mathcal{H}})^{**}(1-z_{K})\to B({\mathcal{H}})$ such that $\Omega(t(1-z_{K}))=t$ for every $t\in{\mathcal{T}}({\mathcal{H}})$
@@
(i) $\Rightarrow$ (iii): By assumption, there is a contraction $a\in{\mathcal{A}}({\mathcal{H}})$ such that $aq=q$ and $\|ap\|<1$ for every closed projection $p\in{\mathcal{T}}({\mathcal{H}})^{**}$ with $pq=0$. Consider the projection $P=\Omega(q(1-z_{K}))$ in $B({\mathcal{H}})$. Since the sequence $(a^{n})$ converges to $q$ in the weak-*** topology of ${\mathcal{T}}({\mathcal{H}})^{**}$ [hay2007, Lemma 3.6], it follows that the sequence $\Omega(a^{n})=a^{n}$ converges to $P$ in the weak-*** topology of $B({\mathcal{H}})$. Then, any vector $f\in{\mathcal{H}}$ in the range of $P$ satisfies $af=f$. Since ${\mathcal{H}}$ consists of analytic functions on the ball, by the identity theorem we conclude that $a=1$ if $P\neq 0$. But $a\neq 1$ since $q$ is ${\mathcal{A}}({\mathcal{H}})$-null. Therefore, $P=0$, and we conclude that $(a^{n})$ converges to 00 in the weak-*** topology of $B({\mathcal{H}})$.
@@
It is trivial that (iv) $\Rightarrow$ (v), while (v)$\Rightarrow$ (iii) follows from Theorem 5.7 under the assumption that there exist non-empty ${\mathcal{M}}({\mathcal{H}})$-totally null sets. ∎
We record some consequences of our arguments so far, using also the facts established in the remark after Theorem 5.6.
@@
Let ${\mathcal{H}}$ be a Hilbert space of analytic functions on ${\mathbb{B}}_{d}$ admitting a reproducing kernel that is unitarily invariant and regular. The ***-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z_{K}\to\mathrm{C}({\mathbb{S}}_{d})^{**}$ restricts to an order preserving bijection between the class of closed projections $q$ characterized in Corollary 5.8 and the class of Bishop peak-interpolation (or equivalently, peak, or ${\mathcal{M}}({\mathcal{H}})$-totally null) subsets $E$ of ${\mathbb{S}}_{d}$.
Using this bijection one may add ‘interpolation set’ to the equivalences in the last line, if nonempty peak sets exist. To close this section, we relate the statements above to the condition that ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$. The key observation is the following. We say that a vector $\xi\in{\mathcal{H}}$ is strictly cyclic for ${\mathcal{A}}({\mathcal{H}})$ if ${\mathcal{A}}({\mathcal{H}})\xi={\mathcal{H}}$.
@@
Let $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ be a closed projection such that ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$. If ${\mathcal{A}}({\mathcal{H}})$ admits no strictly cyclic vector in ${\mathcal{H}}$, then $q\leq z_{K}$.
@@
If we put $P=\Omega(q(1-z_{K}))$ as before, then it suffices to show that $P=0$. As subspaces in ${\mathcal{T}}({\mathcal{H}})^{**}$, we have ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$, which in turn implies ${\mathcal{A}}({\mathcal{H}})q(1-z_{K})={\mathcal{T}}({\mathcal{H}})q(1-z_{K})$. Upon applying $\Omega$, we find ${\mathcal{A}}({\mathcal{H}})P={\mathcal{T}}({\mathcal{H}})P$ as subsets in $B({\mathcal{H}})$. If $P\neq 0$, then there is a unit vector $\xi\in{\mathcal{H}}$ in the range of $P$; let $E\in B({\mathcal{H}})$ denote the corresponding rank-one projection. Since $E\leq P$, we have ${\mathcal{A}}({\mathcal{H}})E={\mathcal{T}}({\mathcal{H}})E$. Let $\eta\in{\mathcal{H}}$ be an arbitrary vector. Choose a rank-one operator $t\in B({\mathcal{H}})$ such that $t\xi=\eta$. Then, $t\in{\mathcal{T}}({\mathcal{H}})$ so there is $a\in{\mathcal{A}}({\mathcal{H}})$ with $aE=tE$. This implies that $a\xi=aE\xi=tE\xi=\eta.$ In other words, ${\mathcal{A}}({\mathcal{H}})\xi={\mathcal{H}}$, so that $\xi$ is strictly cyclic for ${\mathcal{A}}({\mathcal{H}})$. ∎
@@
Let $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ be a closed projection such that ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$. Assume that
@@
    ${\mathcal{A}}({\mathcal{H}})$ admits no strictly cyclic vector in ${\mathcal{H}}$, and
@@
We do not know of a characterization of those spaces ${\mathcal{H}}$ for which ${\mathcal{A}}({\mathcal{H}})$ admits no strictly cyclic vector. Nevertheless, we can give a simple sufficient condition: such vectors cannot exist provided that ${\mathcal{A}}({\mathcal{H}})$ is properly contained in ${\mathcal{M}}({\mathcal{H}})$ (which is the case in many classical examples). The details are as follows. Let $f\in{\mathcal{H}}$ and $b\in{\mathcal{M}}({\mathcal{H}})$. If $f$ is strictly cyclic, then there is $a\in{\mathcal{A}}({\mathcal{H}})$ such that $af=bf$. This equality implies in particular that the analytic functions $af$ and $bf$ agree on the open unit ball. Since $f$ is non-zero, this readily implies $a=b$ using the ‘identity theorem’ from complex analysis. Hence ${\mathcal{A}}({\mathcal{H}})={\mathcal{M}}({\mathcal{H}})$. Some additional partial results on the existence of strictly cyclic vector can be found in [shields1974].
@@ 6. Null projections and the Riesz projection
    $A$ has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
(ii) $\Rightarrow$ (i): Let $\varphi\in A^{\perp}$. Then, $\varphi$ annihilates $A^{\perp\perp}$, and in particular it annihilates $r_{\Delta}A$ since $r_{\Delta}\in A^{\perp\perp}$. This implies in turn that $\varphi(r_{\Delta}\cdot)$ annihilates $A$. ∎
@@
    $A$ has the strong F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
Assume that $\Delta$ is closed in the weak-*** topology of $B^{*}$. Let $\psi$ be a state in $\operatorname{SG}(\Delta)$. Then, there is a type $F_{\sigma}$ projection $q\in B^{**}$ annihilating $\operatorname{AC}(\Delta)$ and satisfying $\psi(q)=1$.
@@
We say that $A$ has the Forelli property in $B$ with respect to $\Delta$ if for every type $F$ projection $q\in B^{**}$ lying in $\operatorname{AC}(\Delta)^{\perp}$, there is a contractive net $(a_{i})$ in $A$ such that $(qa_{i})$ converges to 00 in the weak-*** topology of $B^{**}$, and $(\varphi(1-a_{i}))$ converges to 00 for every $\varphi\in\operatorname{AC}(\Delta)$. If this property only holds for type $F_{\sigma}$ projections, then we say instead that $A$ has the $\sigma$-Forelli property in $B$ with respect to $\Delta$. If this property only holds for type $F_{0}$ projections, then we say instead that $A$ has the finite Forelli property in $B$ with respect to $\Delta$.
@@
    $A$ has the strong F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
    $A$ has the Forelli property in $B$ with respect to $\Delta$.
@@
    $A$ has the $\sigma$-Forelli property in $B$ with respect to $\Delta$.
@@
    $A$ has the finite Forelli property in $B$ with respect to $\Delta$.
@@
    $A$ has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
If we assume in addition that $\Delta$ is weak-*** closed in $B^{*}$, then
@@
(vi) $\Rightarrow$ (vii): Let $q\in B^{**}$ be a type $F$ projection in $\operatorname{AC}(\Delta)^{\perp}$. Thus, there is a collection ${\mathcal{F}}\subset B^{**}$ of closed projections such that $q=\bigvee_{p\in{\mathcal{F}}}p$. Since $q\in\operatorname{AC}(\Delta)^{\perp}$, we see that $q\leq 1-r_{\Delta}$. Hence $p\leq 1-r_{\Delta}$ and so $p\in\operatorname{AC}(\Delta)^{\perp}$ for each $p\in{\mathcal{F}}$. By assumption, for each $p\in{\mathcal{F}}$ there is a projection $r_{p}\in\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$ such that $p\leq r_{p}$. Put $q^{\prime}=\bigvee_{p\in{\mathcal{F}}}r_{p}$, which still lies in $A^{\perp\perp}$. Furthermore, since $r_{p}\in\operatorname{AC}(\Delta)^{\perp}$, we find $r_{p}\leq 1-r_{\Delta}$ for each $p\in F$. Thus, $q^{\prime}\leq 1-r_{\Delta}$ and $q^{\prime}\in\operatorname{AC}(\Delta)^{\perp}$. Now, by virtue of Goldstine’s theorem, because $1-q^{\prime}\in A^{\perp\perp}$, we may find a contractive net $(a_{i})$ in $A$ converging to $1-q^{\prime}$ in the weak-*** topology of $B^{**}$. In particular, because $q\leq q^{\prime}$, this means that $(qa_{i})$ converges to 00 in the weak-*** topology of $B^{**}$. Further, if $\varphi\in\operatorname{AC}(\Delta)$, then since $q^{\prime}\in\operatorname{AC}(\Delta)^{\perp}$ we have $\varphi(q^{\prime})=0$ and
@@
For the rest of the proof, we assume that $\Delta$ is closed in the weak-*** topology of $B^{*}$.
@@
(iv) $\Rightarrow$ (ii): Let $\varphi\in A^{\perp}$ and put $\psi=\varphi((1-r_{\Delta})\cdot)$. We must show that $\psi=0$. Assume towards a contradiction that $\psi\neq 0$. Then, $|\psi|$ lies in $\operatorname{SG}(\Delta)$, and upon scaling if necessary, we may assume $|\psi|$ is a state. By Lemma 6.4, there is a collection ${\mathcal{F}}\subset B^{**}$ of closed projections such that the projection $q=\bigvee_{p\in{\mathcal{F}}}p$ annihilates $\operatorname{AC}(\Delta)$ and satisfies $|\psi|(q)=1$. As before in the proof of (vii), it is readily seen that $p\in\operatorname{AC}(\Delta)^{\perp}$ for every $p\in{\mathcal{F}}$. The assumption then shows that $p$ is $A$-null for each $p\in{\mathcal{F}}$, and thus $q$ is $A$-null. On the other hand, we already know that (iv)$\Rightarrow$(x)$\Leftrightarrow$(xi), so $\psi\in A^{\perp}$ and $|\psi|(q)=0$, which is absurd. ∎
@@
Remark. The reader may notice that there is a notable omission in the long list in the statement of the last theorem, namely that all closed projections in AC$(\Delta)^{\perp}$ are in $A^{\perp\perp}$. This seems to be a very interesting condition. It is equivalent to (iv) in the commutative case, but we are not sure about the general case. Indeed if $B=\mathrm{C}(X)$ for some compact Hausdorff space $X$, then this condition is simply saying that a closed set $E$ in $X$ is a generalized $A$-peak set if each of its closed subsets has $\mu$-measure zero for every probability measure $\mu$ in $\Delta$. That is, a closed set $E\subset X$ is a generalized $A$-peak set if $|\mu|(E)=0$ for every probability measure $\mu$ in $\Delta$. This clearly implies that every closed subset of such a set $E$ is a generalized peak set, which implies by Proposition 2.4 or the discussion above it that $\chi_{E}$ is $A$-null. That is, $E$ is a peak-interpolation set.

### v2

@@ 5.2. Hilbert function spaces and their multipliers
we infer that $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$ is equivalent to $\Theta(q{\mathcal{A}}({\mathcal{H}}))=\Theta(q{\mathcal{T}}({\mathcal{H}}))$, or
$\chi_{E}\,\{\varphi:M_{\varphi}\in{\mathcal{A}}({\mathcal{H}})\}=\chi_{E}\,\mathrm{C}({\mathbb{S}}_{d}).\qed$
We remark here that for a projection $q\leq z_{K}$, the equality in statement (i) above is equivalent to ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$. Indeed, in this case $q$ is central since ${\mathcal{T}}({\mathcal{H}})^{**}z_{K}\cong\mathrm{C}({\mathbb{S}}_{d})^{**}$.
A bounded linear functional $\varphi:{\mathcal{A}}({\mathcal{H}})\to{\mathbb{C}}$ will be said to be ${\mathcal{M}}({\mathcal{H}})$-Henkin if $\varphi$ extends weak-* continuously to ${\mathcal{M}}({\mathcal{H}})$. We note here that this coincides with the notion used in [CT2023Henkin] by Equation (6) above Lemma 5.6 there (or [DH2023, Lemma 2.1]). This will be used implicitly below, whenever results from [CT2023Henkin] regarding Henkin functionals are invoked.
@@
$I_{\mu}(f)=\int_{{\mathbb{S}}_{d}}fd\mu,\quad f\in\mathrm{C}({\mathbb{S}}_{d}).$
A closed projection $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ is ${\mathcal{M}}({\mathcal{H}})$-totally null if $|\varphi|(q)=0$ for every functional $\varphi\in{\mathcal{T}}({\mathcal{H}})^{*}$ such that $\varphi|_{{\mathcal{A}}({\mathcal{H}})}$ is ${\mathcal{M}}({\mathcal{H}})$-Henkin. A Borel subset $E\subset{\mathbb{S}}_{d}$ is said to be ${\mathcal{M}}({\mathcal{H}})$-totally null if $|\mu|(E)=0$ for every ${\mathcal{M}}({\mathcal{H}})$-Henkin measure. We clarify the relation between these notions in the next result. In what follows, we will repeatedly use the fact that $\varphi(\cdot(1-z_{K}))|_{{\mathcal{A}}({\mathcal{H}})}$ is an ${\mathcal{M}}({\mathcal{H}})$-Henkin functional for every $\varphi\in{\mathcal{T}}({\mathcal{H}})^{*}$ (see [CT2023Henkin, Lemma 5.5]).
@@
Since $q$ is closed and ${\mathcal{T}}({\mathcal{H}})$ is finitely generated by the monomials, hence is separable, we may choose a sequence of positive contractions $(t_{n})$ in ${\mathcal{T}}({\mathcal{H}})$ decreasing to $q$ in the weak-* topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. In particular, we see that the bounded sequence of continuous functions $((\tau\circ\gamma)(t_{n}))$ converges pointwise on ${\mathbb{S}}_{d}$ to $\chi_{E}$. By dominated convergence and using that $(\tau\circ\nu)^{**}(z_{K})=1$, we then find
@@ 5.3. The Davidson–Hartz theorems
The previous result has bearing on the topic of peak-interpolation. Indeed, it allows us to provide alternative proofs (and even some refinements) of some of the results from [DH2023] via the noncommutative peak-interpolation theory surveyed in [blecher2013]. For the most part, the statement below is a commutative one. That our approach takes us through the abstract noncommutative theory is a novel feature that we feel has potential for further future applications.
@@
Let $g\in\mathrm{C}(E)$ with $\|g\|=1$. Let $k_{0}\in\mathrm{C}({\mathbb{S}}_{d})$ be a strictly positive function such that $k_{0}|_{E}=1$ and $0<k<1$ on ${\mathbb{S}}_{d}\setminus E$. Let $k\in\mathrm{C}({\mathbb{S}}_{d})$ be strictly positive such that $|g|\leq k$ on $E$ and put $k_{1}=\min\{k,k_{0}\}$. By Tietze’s extension theorem, there is $g_{0}\in\mathrm{C}({\mathbb{S}}_{d})$ such that $g_{0}|_{E}=g$ and $\|g_{0}\|=1$. Recall that $\tau\circ\gamma:{\mathcal{T}}({\mathcal{H}})\to\mathrm{C}({\mathbb{S}}_{d})$ is a surjective *-homomorphism. Thus, there is $b\in{\mathcal{T}}({\mathcal{H}})$ with $\|b\|=1$ and a positive element $d\in{\mathcal{T}}({\mathcal{H}})$ with $\|d\|=\|k_{1}\|$ such that $(\tau\circ\gamma)(b)=g_{0}$ and and $(\tau\circ\gamma)(d)=k_{1}$.
@@
Using that $E$ is an ${\mathcal{A}}({\mathcal{H}})$-interpolation set, we find $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$ by Lemma 5.4. Recall that ${\mathcal{T}}({\mathcal{H}})^{**}z_{K}\cong\mathrm{C}(\mathbb{S}_{d})^{**}$ and $q\leq z_{K}$, so that $q$ is central. Since ${\mathcal{T}}({\mathcal{H}})$ is nuclear, by Proposition 4.8 we infer that there is $\kappa>0$ such that $q$ is a complete $({\mathcal{A}}({\mathcal{H}}),\kappa)$-interpolation projection.
@@
The following result closes the circle of results about peak and null sets and projections for ${\mathcal{A}}({\mathcal{H}})$. Recall that by [CT2023Henkin, Lemma 5.5], there is a weak-* homeomorphic *-isomorphism $\Omega:{\mathcal{T}}({\mathcal{H}})^{**}(1-z_{K})\to B({\mathcal{H}})$ such that $\Omega(t(1-z_{K}))=t$ for every $t\in{\mathcal{T}}({\mathcal{H}})$.
@@
(i) $\Rightarrow$ (iii): By assumption, there is a contraction $a\in{\mathcal{A}}({\mathcal{H}})$ such that $aq=q$ and $\|ap\|<1$ for every closed projection $p\in{\mathcal{T}}({\mathcal{H}})^{**}$ with $pq=0$. Consider the projection $P=\Omega(q(1-z_{K}))$ in $B({\mathcal{H}})$. Since the sequence $(a^{n})$ converges to $q$ in the weak-* topology of ${\mathcal{T}}({\mathcal{H}})^{**}$ [hay2007, Lemma 3.6], it follows that the sequence $\Omega(a^{n}(1-z_{K}))=a^{n}$ converges to $P$ in the weak-* topology of $B({\mathcal{H}})$. Then, any vector $f\in{\mathcal{H}}$ in the range of $P$ satisfies $af=f$. Since $\mathcal{H}$} consists of analytic functions on the ball, by the identity theorem we conclude that $a=1$ if $P\neq 0$. But $a\neq 1$ since $q$ is ${\mathcal{A}}({\mathcal{H}})$-null. Therefore, $P=0$, and we conclude that $(a^{n})$ converges to 0 in the weak-* topology of $B({\mathcal{H}})$.
@@
[…]
We record some consequences of our arguments so far, using also the facts established in the remark after Theorem 5.6. We reiterate that we are recovering results from [DH2023] here using our techniques.
@@
Let $\mathcal{H}$} be a Hilbert space of analytic functions on ${\mathbb{B}}_{d}$ admitting a reproducing kernel that is unitarily invariant and regular. The *-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z_{K}\to\mathrm{C}({\mathbb{S}}_{d})^{**}$ restricts to an order preserving bijection between the class of closed projections $q$ characterized in Corollary 5.8 and the class of Bishop peak-interpolation (or equivalently, peak, or ${\mathcal{M}}({\mathcal{H}})$-totally null) subsets $E$ of ${\mathbb{S}}_{d}$.
Using this bijection one may add ‘interpolation set’ to the equivalences in the last line, if nonempty peak sets exist. To close this section, we relate the statements above to the condition that $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$. The key observation is the following. We say that a vector $\xi\in{\mathcal{H}}$ is strictly *-cyclic for ${\mathcal{A}}({\mathcal{H}})$ if ${\mathcal{A}}({\mathcal{H}})^{*}\xi={\mathcal{H}}$.
@@
Let $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ be a closed projection such that $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$. If ${\mathcal{A}}({\mathcal{H}})$ admits no strictly *-cyclic vector in $\mathcal{H}$}, then $q\leq z_{K}$.
@@
If we put $P=\Omega(q(1-z_{K}))$ as before, then it suffices to show that $P=0$. As subspaces in ${\mathcal{T}}({\mathcal{H}})^{**}$, we have ${\mathcal{A}}({\mathcal{H}})^{*}q={\mathcal{T}}({\mathcal{H}})q$, which in turn implies ${\mathcal{A}}({\mathcal{H}})^{*}q(1-z_{K})={\mathcal{T}}({\mathcal{H}})q(1-z_{K})$. Upon applying $\Omega$, we find ${\mathcal{A}}({\mathcal{H}})^{*}P={\mathcal{T}}({\mathcal{H}})P$ as subsets in $B({\mathcal{H}})$. If $P\neq 0$, then there is a unit vector $\xi\in{\mathcal{H}}$ in the range of $P$; let $E\in B({\mathcal{H}})$ denote the corresponding rank-one projection. Since $E\leq P$, we have ${\mathcal{A}}({\mathcal{H}})^{*}E={\mathcal{T}}({\mathcal{H}})E$. Let $\eta\in{\mathcal{H}}$ be an arbitrary vector. Choose a rank-one operator $t\in B({\mathcal{H}})$ such that $t\xi=\eta$. Then, $t\in K({\mathcal{H}})\subset{\mathcal{T}}({\mathcal{H}})$ so there is $a\in{\mathcal{A}}({\mathcal{H}})$ with $a^{*}E=tE$. This implies that $a^{*}\xi=a^{*}E\xi=tE\xi=\eta.$ In other words, ${\mathcal{A}}({\mathcal{H}})^{*}\xi={\mathcal{H}}$, so that $\xi$ is strictly *-cyclic for ${\mathcal{A}}({\mathcal{H}})$. ∎
@@
Let $q\in{\mathcal{T}}({\mathcal{H}})^{**}$ be a closed projection such that $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$. Assume that
@@
    ${\mathcal{A}}({\mathcal{H}})$ admits no strictly *-cyclic vector in $\mathcal{H}$}, and
@@
We do not know of a characterization of those spaces $\mathcal{H}$} for which ${\mathcal{A}}({\mathcal{H}})$ admits no strictly *-cyclic vector. Nevertheless, we can give a simple sufficient condition: such vectors cannot exist provided that ${\mathcal{A}}({\mathcal{H}})$ is properly contained in ${\mathcal{M}}({\mathcal{H}})$ (which is the case in many classical examples). Indeed, this follows from the following general fact.
If $A\subset B(H)$ is a commutative subalgebra properly contained in its commutant $A^{\prime}$, then $A$ admits no strictly *-cyclic vector. To see this, assume that $\xi\in H$ is strictly *-cyclic for $A$, and choose $b\in A^{\prime}$ with $b\notin A$. Then, there is $a\in A$ such that $a^{*}\xi=b^{*}\xi$. Using that $A$ is commutative, this further implies $a^{*}c^{*}\xi=b^{*}c^{*}\xi$ for every $c\in A$, and in turn we must have $a^{*}=b^{*}$ or $a=b$, a contradiction.
Some additional partial results on the existence of strictly cyclic vector can be found in [shields1974].
@@ 6. Null projections and the Riesz projection
    AA has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
(ii) $\Rightarrow$ (i): Let $\varphi\in A^{\perp}$. Then, $\varphi$ annihilates $A^{\perp\perp}$, and in particular it annihilates $r_{\Delta}A$ since $r_{\Delta}\in A^{\perp\perp}=\bar{A}^{w*}$. This implies in turn that $\varphi(r_{\Delta}\cdot)$ annihilates $A$. ∎
@@
    AA has the strong F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
Assume that $\Delta$ is closed in the weak-* topology of $B^{*}$. Let $\psi$ be a state in $\operatorname{SG}(\Delta)$. Then, there is a type $F_{\sigma}$ projection $q\in B^{**}$ annihilating $\operatorname{AC}(\Delta)$ and satisfying $\psi(q)=1$.
@@
We say that $A$ has the Forelli property in $B$ with respect to $\Delta$ if for every type $F$ projection $q\in B^{**}$ lying in $\operatorname{AC}(\Delta)^{\perp}$, there is a contractive net $(a_{i})$ in $A$ such that $(qa_{i})$ converges to 0 in the weak-* topology of $B^{**}$, and $(\varphi(1-a_{i}))$ converges to 0 for every $\varphi\in\operatorname{AC}(\Delta)$. If this property only holds for type $F_{\sigma}$ projections, then we say instead that $A$ has the $\sigma$-Forelli property in $B$ with respect to $\Delta$. If this property only holds for type $F_{0}$ projections, then we say instead that $A$ has the finite Forelli property in $B$ with respect to $\Delta$.
@@
    AA has the strong F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
    AA has the Forelli property in $B$ with respect to $\Delta$.
@@
    AA has the $\sigma$-Forelli property in $B$ with respect to $\Delta$.
@@
    AA has the finite Forelli property in $B$ with respect to $\Delta$.
@@
    AA has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
If we assume in addition that $\Delta$ is weak-* closed in $B^{*}$, then
@@
(vi) $\Rightarrow$ (vii): Let $q\in B^{**}$ be a type $F$ projection in $\operatorname{AC}(\Delta)^{\perp}$. Thus, there is a collection ${\mathcal{F}}\subset B^{**}$ of closed projections such that $q=\bigvee_{p\in{\mathcal{F}}}p$. Since $q\in\operatorname{AC}(\Delta)^{\perp}$, we see that $q\leq 1-r_{\Delta}$. Hence $p\leq 1-r_{\Delta}$ and so $p\in\operatorname{AC}(\Delta)^{\perp}$ for each $p\in{\mathcal{F}}$. By assumption, for each $p\in{\mathcal{F}}$ there is a projection $r_{p}\in\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$ such that $p\leq r_{p}$. Put $q^{\prime}=\bigvee_{p\in{\mathcal{F}}}r_{p}$, which still lies in $A^{\perp\perp}$. Furthermore, since $r_{p}\in\operatorname{AC}(\Delta)^{\perp}$, we find $r_{p}\leq 1-r_{\Delta}$ for each $p\in F$. Thus, $q^{\prime}\leq 1-r_{\Delta}$ and $q^{\prime}\in\operatorname{AC}(\Delta)^{\perp}$. Now, by virtue of Goldstine’s lemma, because $1-q^{\prime}\in A^{\perp\perp}$, we may find a contractive net $(a_{i})$ in $A$ converging to $1-q^{\prime}$ in the weak-* topology of $B^{**}$. In particular, because $q\leq q^{\prime}$, this means that $(qa_{i})$ converges to 0 in the weak-* topology of $B^{**}$. Further, if $\varphi\in\operatorname{AC}(\Delta)$, then since $q^{\prime}\in\operatorname{AC}(\Delta)^{\perp}$ we have $\varphi(q^{\prime})=0$ and
@@
For the rest of the proof, we assume that $\Delta$ is closed in the weak-* topology of $B^{*}$.
@@
[…]
@@
[…]